L has y-intercept (0,2) and is perpendicular to the line with equation 2x- 3y= 6.
Write the equation of the line passing through each of the given pair points. Whats the result in solpe-intercept form ?

Answer by venugopalramana(3286)

L has y-intercept (0,2)...(X1,Y1)
HENCE EQN .IS
Y-Y1=M(X-X1)...WHERE M IS SLOPE
Y-2=M(X-0)=MX
and is perpendicular to the line with equation 2x- 3y= 6.
SLOPE OF THIS LINE......3Y=2X-6.....OR.....Y=2X/3-2.......
SLOPE =2/3
SLOPE OF PERPENDICULAR =M=-3/2
EQN.OF L IS
Y-2=-3X/2.........2Y-4=-3X.....3X+2Y-4=0
Write the equation of the line passing through each of the given pair points.
WHAT POINTS?
Whats the result in solpe-intercept form ?
Y=-1.5X+2